### v1

@@ 3.5 Mode Gruneisen parameters
The mode Gruneisen parameters are obtained from the phonons and the cubic force constants. They represent the change in the phonon frequencies under applied strain. Even though the strain is usually considered to be iosotropic compression or dilation, one can also define the response to an arbitrary strain. By definition, we have $\gamma_{\lambda,\mu\nu}=-\frac{dln\,\omega_{\lambda}}{3d\epsilon_{\mu\nu}}$. But usually the volume derivative, which corresponds to an isotropic strain is considered: $\gamma_{\lambda}=-\frac{dln\,\omega_{\lambda}}{dln\,V}$ where $V$ is the unit cell volume. Applying a uniform hydrostatic strain $\epsilon_{\alpha\beta}=\delta_{\alpha\beta}\,dV/3V=\delta_{\alpha\beta}\,dln\,V/3$. Similar to the group velocity, we start with the square of the frequency written as the product of the dynamical matrix by eigenvectors: $\omega^{2}=e.D.e$ and consider the change in $D$ under strain: $dD/d\eta=\sum_{R}d\Phi/d\eta\,e^{ik.R}$, and use $d\Phi/d\eta=\Psi(R+\tau+du^{0}/d\eta)$. Note the latter term, which is non-zero for low-symmetry structures having their own internal relaxation under strain $u^{0}(\eta)$, is often neglected in the literature. In the previous section, we derived this equation $u^{0}(\eta)=-\Gamma(\Pi+Q\eta)$ for $u^{0}$ in the harmonic approximation. Therefore $du^{0}/d\eta=-\Gamma Q$. Using the Hellmann-Feynman theorem applied to the derivative of the dynamical matrix, and substituting the above relation for the strain derivative of the harmonic force constants, we finally find:
$\gamma_{\lambda}(k)=-\frac{1}{6\omega_{\lambda}^{2}(k)}\sum_{\tau,R_{1}\tau_{1},R_{2}\tau_{2}}\frac{\Psi_{\tau,R_{1}\tau_{1},R_{2}\tau_{2}}^{\alpha\alpha_{1}\alpha_{2}}}{\sqrt{m_{\tau}m_{\tau_{1}}}}\,e_{\tau\alpha,\lambda}(k)^{*}\,e_{\tau_{1}\alpha_{1},\lambda}(k)\,e^{ik.(R_{1}+\tau_{1}-\tau)}\,\left[(R_{2}+\tau_{2})^{\alpha_{2}}-\sum_{\mu=1,2,3}\Gamma_{\tau_{2}\alpha_{2}},.Q_{.,\mu\mu}\right]$ (24)
@@ 3.6 Thermodynamic properties within the quasi-harmonic approximation (QHA)
$F^{\rm QHA}(T)\approx F(V(T))=E_{0}(V(T))+k_{B}T\sum_{\lambda}{\rm ln}\,[2\,{\rm sinh}(\hbar\omega(V(T))/2k_{B}T)]$ (25)
@@
For the zero-temperature part of the free or total energy $E_{0}$ one can either assume $E_{0}(V)=E_{0}(V_{0})+\frac{1}{2V_{0}}B(V-V_{0})^{2}$, with $B$ being the bulk modulus, or just use the Taylor expansion 19 in powers of $\eta$ and replace volume derivatives by strain derivatives: $dlnV=dV/V_{0}=3d\epsilon$, so that the zero-temperature equation of state can be written as:
@@
The third derivative can be easily obtained only if $u^{0}=0$ due to symmetry; otherwise the expression becomes complicated, due to the fact that equation 13 has to be extended to third-order and equation 18 would not be linear in $\eta$ anymore, and needs to be solved for $\eta$ and $\eta$ be substituted back into 13 before taking the derivatives. Only in the case where $u^{0}=0$ or $du^{0}/d\eta=0$, do we have
$\frac{d^{3}E_{0}}{d\epsilon^{3}}(u^{0}=0;du^{0}/d\eta=0)=\frac{1}{6N}\sum_{R_{i}\tau_{i}\alpha_{i}}\Psi_{R_{1}\tau_{1},R_{2}\tau_{2},R_{3}\tau_{3}}^{\alpha_{1}\alpha_{2}\alpha_{3}}(R_{1}+\tau_{1})^{\alpha_{1}}(R_{2}+\tau_{2})^{\alpha_{2}}(R_{3}+\tau_{3})^{\alpha_{3}}$
In this case, we can also deduce the dimensionless parameter, so-called $B^{\prime}_{0}=dB/dP=(V_{0}d^{3}E/dV^{3})/(d^{2}E/dV^{2})$.
To find the volume dependence of the phonon frequencies, we use $d\omega_{\lambda}/dV=-\gamma_{\lambda}\omega_{\lambda}/V$, and can finally find the equilibrium volume $V_{eq}(T)$ by minimizing $F^{\rm QHA}$ with respect to the strain or the volume.
Since the bulk modulus itself is a function of $T$, or, within the QHA, of the volume, we need to consider this dependence in the first term of the free energy when performing volume-minimization. The simplest volume dependence of the bulk modulus is a linear one. So we take $B(V)=B(V_{0})(1-B^{\prime}_{0}(V-V_{0})/V_{0})$ In FOCEX, to include a temperature-dependence in the bulk modulus, we therefore use $B(V)=B(V_{0})(1-3\eta_{\rm eq}(T)B^{\prime}_{0})$. Note that only for materials with no internal relaxations $u^{0}$ can we calculate and use $B^{\prime}_{0}$.
@@ 3.8 Brief algorithm
2) Force-displacement supercell data must be stored in files named POSCAR1, POSCAR2, ... ; FORCEDISP1, FORCEDISP2, .... The FOCEX code checks consistency between the atomic structures of the supercell and the primitive cell described in structure.params. Force data may come from supercells of different shapes. Data for each shape is stored in a different input POSCARi, ... ; FORCEDISPi, ... i=1,2,… . Typically one large supercell can be used to extract the harmonic FCs, which can be of long range, and a smaller supercell can be used for large-amplitucde anharmonic displacements. The files POSCARi contain the equilibrium or reference coordinates of the atoms in the supercell number i, in the POSCAR format defined in VASP.
@@ 3.9 Input files
3) The information regarding the reference supercells are read from files POSCAR1, POSCAR2, ... written in the same format as VASP POSCAR file. The corresponding cartesian displacements (in Ang) and forces (in eV/Ang) are read from files FORCEDISP1, FORCEDISP2, .... Each FORCEDISPi file contains all the force-displacement snapshots corresponding to the $i^{th}$ supercell defined in POSCARi. Each snapshot is separated from the next one by two lines, the first of which should contain the word "POSITION" and the second one contains the total energy of that snapshot. The N following lines contain the cartesian coordinates of the N atoms followed by the forces acting on them (6 columns in total).
4) The file kpbs.params contains the list of kpoints to use for the phonon band structure calculations. After the force constants are extracted, the FOCEX code also computes the harmonic properties, namely the phonon band structure, elastic and thermodynamic properties.
@@ 3.10 Output files
4) log-<extra>.dat contains the log of the run, where the parameters "<extra>" label the type of the run, as described in the file ios.f90.
@@
9) ibz_dos.dat contains the Gaussian-smeared DOS from kpoints generated in the IBZ, while fbz_dos.dat contains the DOS generated using the tetrahedron method.
@@ 3.11 Utilities
Some utilities (read_outcar.f90 and read_qe.f90) are also provided to convert the log or output files from a single snapshot of VASP (OUTCAR) or Quantum Espresso to FORCEDISP format. A shell script xtract.sh then calls these executables, and appends the output from each snapshot into the file FORCEDISP1. In addition, files for plotting the outputs using the Gnuplot software are also provided. all.plt or bvel.plt plot the phonon band structure, Gruneisen parameters DOS, and the group velocities. KTICS.BS should be imported in this file. fcs.plt plots the harmonic force constants versus pair distance and visualizes them in the Wigner-Seitz cell associated with the supercell. qha.plt plots the results from the QHA calculations: total energy, free energy, strain, volume, heat capacity, entropy and bulk modulus versus temperature. kp.plt visualizes the full zone, irreducible zone and band structure kpoints in 3d reciprocal space.
@@ 4 ANFOMOD
To update atomic positions, the standard velocity Verlet Ohno et al. (2018) algorithm is used. If canonical ensemble needs to be simulated, a Langevin thermostat, in which the damping rate can be adjusted, has been adopted. A good initial guess for the latter would be $1/(1000dt)$ where $dt$ is the MD timestep.
@@ 6 Installation
The ALADYN software package is available on GitHub under GPL-3.0 license at https://github.com/KeivanS/Anharmonic-lattice-dynamics, where each code is stored in the corresponding subdirectory. A Makefile template for each code has been provided. After cloning the repository, the user can type make to generate all needed executables, or alternatively go into each subdirectory, edit the Makefile according to their system and needs, and run make. The codes can be compiled using the GNU Fortran Compiler gfortran or the Intel Fortran Compiler ifort, and are tested to be compatible with gfortran versions 7-11 and ifort versions 18-19. (Newer and/or older versions are likely to be compatible as well.) The only external library is the zhegv diagonalization routine from LAPACK. A successful compilation results in the creation of the ALADYN_BIN directory containing all the executables.
In a typical high-performance computing cluster environment, one should be able to compile the codes using commands similar to the following. For example, to compile FOCEX with GCC (assuming gfortran is the installed fortran compiler):
@@ 7 Example of Ge, and some tips
Here we show the example of Germanium as a system ran with FOCEX. We must emphasize that the force-displacements inputs stored in FORCEDISP1 are up to the user. For a choice of the supercell and magnitude of displacements, the best set of FCs will be extracted. But the results will change with a change in the supercell, magnitude of displacements, choice of the exchange-correlation potential, lattice parameter, chosen rank and chosen range of FCs. So it is up to the user to make sure the obtained results are converged. As a rule of thumb, a supercell of 100 atoms or more is usually enough to provide a very reasonable set of force constants. To get accurate harmonic force constants, displacements need to be small enough. Typically displacements on the order of or less than 0.05 Å  are reasonable. While many codes require moving one or two atoms at the most, in FOCEX all atoms maybe simultaneously moved, providing more information on FCs for the same data size. A reasonable scheme would also be to move all atoms as a superposition of the normal modes sampled from the canonical ensemble, if effective force constants at higher temperatures need to the be extracted. If normal modes are not known, as is usually the case, one can still adopt normal modes generated with an arbitrary but cheap force field. This is done in the utility file sc_snaps.x which generates supercell snapshots of a required number if a temperature of interest is given. In principle, one or two snapshots in which all atoms are moved maybe sufficient to generate harmonic force constants. If only harmonic FCs are required, it is best to generate snapshots in which only one inequivalent atom is moved. More snapshots will be needed if anharmonic FCs are also required, typically on the order of 20, depending on their rank and chosen range. If the range of the required anharmonic FCs is increased, their number is also exponentially increased, and therefore more data (snapshots) would be needed to solve for all the unknowns. Checks must be performed to make sure the generated FCs don’t change when the range is increased; just the furthest ones are added and are smaller than the existing ones. Another check for convergence is that the physical quantities of interest such as the phonon dispersion and the Gruneisen parameters do not appreciably change as ranges are increased. Largest (infinite norm) and average (first norm) error, as well as the percentage deviation (second norm) on fitting are written at the end of the log file.
@@
For Ge, we have used a supercell of 216 atoms with 36 snapshots in which only one or two atoms have been moved. Forces were calculated using VASPKresse and Hafner (1993, 1994); Kresse and Furthmüller (1996a, b) with the GGA-PBEsol functional, at cutoff energy of 225 eV and a Gamma-centered 7x7x7 kpoint mesh within the 3x3x3 supercell. The lattice constant of $a=5.702$was found using 14x14x14 kmesh for the primitive cell optimization calculations. This is only 0.7% larger than the experimental value of $5.66$. All harmonic FCs within the supercell were included in the fitting as shown in Fig. 2. Cubic force constants were included up to the 6th nearest neighbors. The results on the phonon dispersion and Gruneisen parameters are displayed in Fig. 3. The slight discrepancy between the computed and experimentally measuredNilsson and Nelin (1972) frequencies is most likely due to the larger PBE lattice constant of 5.7 Å, which is slightly larger than the experimental value of 5.66 Å. This is a 2.1% larger volume. Therefore for a Gruneisen parameter of +1 for instance, a 2.1% reduction in the phonon frequency should be expected. The Gruneisen parameters in Fig.3 show a very decent agreement with the values of Olego and Cardona Olego and Cardona (1982).
@@
Figure 2: Output of the fc.plt: On the left are the fitted Ge harmonic force constants as a function of pair distance. Bottom right: To visualize included force constants, we show the Wigner-Seitz (WS) cell of the supercell (in blue) centered on an atom, the grid of primitive translation vectors (green spheres), and selected force constants (purple lines) connecting the central atom to its neighbors within the supercell. This can be visualized for every atom in the primitive cell. Top right: Wigner-Seitz cell of the Brillouin zone (purple dots).
@@
Figure 3: Mode Gruneisen parameters in the irreducible Brillouin zone (left), phonon density of states (center), and dispersion of Ge (right) calculated by FOCEX. Large red dots are experimental data. Color bar shows the group velocity magnitude in km/s
@@
This work (LD model) 121 69 54.0 86.4 73 105 54 41 0.36 0.29
Experiment (Ioffe) 126 44 67.7 71.3 103 41 0.26
Experiment (McSkimin) 129 48.3 67.25 77 102.5 67.1 0.26
Materials Project (DFT) 102 36 56 58 83.3 47 0.25
Table 1: Elastic constants of Ge crystal in GPa. Experimental data from McSkiminMcSkimin (1953) is at T=298K, and those of Ioffe Institute Database (????) are at 300K. The reported value of 0.19 on MP is not consistent with $-S_{12}/S_{11}=0.25$, but we are reporting the latter. For the Young and shear moduli and the Poisson ratio, we are reporting them for both deformations along [100] and the average value denoted by the subscript iso.
Group velocities are also compared to the data in Ioffe siteDatabase (????). Results, which are presented in table 2, show relatively good agreement. It is worth noting that there is a simple relationship between group velocities near the zone center (sound speeds) and elastic constants in a cubic crystal. All 3 elastic constants can be extracted from the 3 speeds of sound along [110]:
$\rho\,v_{L}^{2}(100)$ $=C_{11}\,\,$ $\rho\,v_{T1}^{2}(100)$ =\,= $\rho\,v_{T2}^{2}(100)$ $=C_{44}$ (32)
$\rho\,v_{L}^{2}(111)$ $=\left(\frac{C_{11}+2C_{12}+4C_{44}}{3}\right)\,\,$ $\rho\,v_{T1}^{2}(111)$ $=$ $\rho\,v_{T2}^{2}(111)$ $=\left(\frac{C_{11}-C_{12}+C_{44}}{3}\right)$ (33)
$\rho\,v_{L}^{2}(110)$ $=\left(\frac{C_{11}+C_{12}}{2}+C_{44}\right)\,\,$ $\rho\,v_{T1}^{2}(110)$ $=C_{44}\,\,$ $\rho\,v_{T2}^{2}(110)$ $=\left(\frac{C_{11}-C_{12}}{2}\right)$ (34)
@@
Table 2: Group velocities (km/s) of the three acoustic branches compared to experimental data from McSkiminMcSkimin (1953) and Ioffe InstituteDatabase (????) taken at room temperature. The largest number corresponds to the longitudinal mode. Along [100] and [111] the two transverse acoustic branches are degenerate.
Finally, we must add that this code has been used many times and its results validated in the past 15 years or so. The results on force constants or phonons and lattice thermal conductivity have been published as follows: SiliconEsfarjani et al. (2011); Minnich et al. (2011); Tian et al. (2011); Scott et al. (2018), half-heuslersShiomi et al. (2011) FeSb2Liao et al. (2014), grapheneMingo et al. (2010); Lee et al. (2015), GaAsMaznev et al. (2013), PbSe and PbTeShiga et al. (2012); Tian et al. (2012), Bi and SbLee et al. (2015).
@@
The Anharmonic LAttice DYNamics (ALADYN) suites of codes may be accessed and downloaded from
@@
https://aladyn.readthedocs.io/en/latest/index.html

### v2

@@ 3.5 Mode Gruneisen parameters
The mode Gruneisen parameters are obtained from the phonons and the cubic force constants. They represent the change in the phonon frequencies under applied strain. Even though the strain is usually considered to be isotropic compression or dilation, one can also define the response to an arbitrary strain. By definition, we have $\gamma_{\lambda,\mu\nu}=-\frac{dln\,\omega_{\lambda}}{3d\epsilon_{\mu\nu}}$. But usually the volume derivative, which corresponds to an isotropic strain is considered: $\gamma_{\lambda}=-\frac{dln\,\omega_{\lambda}}{dln\,V}$ where $V$ is the unit cell volume. Applying a uniform hydrostatic strain $\epsilon_{\alpha\beta}=\delta_{\alpha\beta}\,dV/3V=\delta_{\alpha\beta}\,dln\,V/3$. Similar to the group velocity, we start with the square of the frequency written as the product of the dynamical matrix by eigenvectors: $\omega^{2}=e.D.e$ and consider the change in $D$ under strain: $dD/d\eta=\sum_{R}d\Phi/d\eta\,e^{ik.R}$, and use $d\Phi/d\eta=\Psi(R+\tau+du^{0}/d\eta)$. Note the latter term, which is non-zero for low-symmetry structures having their own internal relaxation under strain $u^{0}(\eta)$, is often neglected in the literature. In the previous section, we derived this equation $u^{0}(\eta)=-\Gamma(\Pi+Q\eta)$ for $u^{0}$ in the harmonic approximation. Therefore $du^{0}/d\eta=-\Gamma Q$. Using the Hellmann-Feynman theorem applied to the derivative of the dynamical matrix, and substituting the above relation for the strain derivative of the harmonic force constants, we finally find:
$\gamma_{\lambda}(k)=-\frac{1}{6\omega_{\lambda}^{2}(k)}\sum_{\tau,R_{1}\tau_{1},R_{2}\tau_{2}}\frac{\Psi_{\tau,R_{1}\tau_{1},R_{2}\tau_{2}}^{\alpha\alpha_{1}\alpha_{2}}}{\sqrt{m_{\tau}m_{\tau_{1}}}}\,e_{\tau\alpha,\lambda}(k)^{*}\,e_{\tau_{1}\alpha_{1},\lambda}(k)\,e^{ik.(R_{1}+\tau_{1}-\tau)}\\ \times\,\left[(R_{2}+\tau_{2})^{\alpha_{2}}-\sum_{\mu=1,2,3}\Gamma_{\tau_{2}\alpha_{2}},.Q_{.,\mu\mu}\right]$ (22)
@@ 3.6 Thermodynamic properties within the quasi-harmonic approximation (QHA)
${\cal F}^{\rm QHA}(T)\approx{\cal F}(V(T))=E_{0}(V(T))+k_{B}T\sum_{\lambda}{\rm ln}\,[2\,{\rm sinh}(\hbar\omega(V(T))/2k_{B}T)]$ (23)
@@
For the zero-temperature part of the free or total energy $E_{0}$ one can either assume $E_{0}(V)=E_{0}(V_{0})+\frac{1}{2V_{0}}B(V-V_{0})^{2}$, with $B$ being the bulk modulus, or just use the Taylor expansion 17 in powers of $\eta$ and replace volume derivatives by strain derivatives: $dlnV=dV/V_{0}=3d\epsilon$, so that the zero-temperature equation of state can be written as:
@@
The third derivative can be easily obtained only if $u^{0}=0$ due to symmetry; otherwise the expression becomes complicated, due to the fact that equation 15 has to be extended to third-order and equation 16 would not be linear in $\eta$ anymore, and needs to be solved for $\eta$ and $\eta$ be substituted back into 15 before taking the derivatives. Only in the case where $u^{0}=0$ or $du^{0}/d\eta=0$, do we have
$\frac{d^{3}E_{0}}{d\epsilon^{3}}(u^{0}=0;\frac{du^{0}}{d\eta}=0)=\\ \frac{1}{6N}\sum_{R_{i}\tau_{i}\alpha_{i}}\Psi_{R_{1}\tau_{1},R_{2}\tau_{2},R_{3}\tau_{3}}^{\alpha_{1}\alpha_{2}\alpha_{3}}(R_{1}+\tau_{1})^{\alpha_{1}}(R_{2}+\tau_{2})^{\alpha_{2}}(R_{3}+\tau_{3})^{\alpha_{3}}$ (26)
In this case, we can also deduce the dimensionless parameter
$B^{\prime}_{0}=\left(\frac{dB}{dP}\right)_{T=0}=V_{0}\left(\frac{d^{3}E}{dV^{3}}\right)/\left(\frac{d^{2}E}{dV^{2}}\right)=\frac{1}{27V_{0}B}\left(\frac{d^{3}E_{0}}{d\epsilon^{3}}\right)$
The volume dependence of the phonon frequencies is given by this formula: $d\omega_{\lambda}/dV=-\gamma_{\lambda}\omega_{\lambda}/V$ We can finally find the equilibrium volume $V_{eq}(T)$ by minimizing $F^{\rm QHA}$ with respect to the strain or the volume.
Since the bulk modulus itself is a function of $T$, or, within the QHA, of the volume, we need to consider this dependence in the first term of the free energy when performing volume-minimization. The simplest volume dependence of the bulk modulus is a linear one. So we take $B(V)=B(V_{0})(1+B^{\prime}_{0}(V-V_{0})/V_{0})$ In FOCEX, to include a temperature-dependence in the bulk modulus, we therefore use $B(V)\approx B(V_{0})(1+3\eta_{\rm eq}(T)B^{\prime}_{0})$. Note that only for materials with no internal relaxations $u^{0}$ can we easily calculate and use $B^{\prime}_{0}$.
3.7 Beyond QHA
If one has access to the mode Gruneisen parameters, the QHA is the simplest way to describe thermal expansion in a crystal. To go beyond QHA, one needs to consider the more accurate anharmonic form of the Free energy. Its volume derivative gives the pressure (mechanical equation of state). Solving the zero pressure equation for the volume then provides the equilibrium volume and thus the thermal expansion. Wallace has a discussion of the free energy of an anharmonic system in [25]. We also provide a variational formulation of the free energy in [27]. In the latter, the free energy is a sum of an effective harmonic free energy, where the potential energy is defined by $K$ as in eq. 14, and an anharmonic correction as defined by $\langle V^{LD}-K_{ij}y_{i}y_{j}/2\rangle$. This can be a starting point for taking volume or strain derivatives and solve $P(V_{eq})=0$. More on this will be published in our upcoming paper on SCOP8. This discussion is limited to insulators when the electronic contribution to thermal expansion is neglected. Allen [29] provides a more extensive discussion of this issue including electronic excitation effects also valid for metals.
@@ 3.9 Brief algorithm
2) Force-displacement supercell data must be stored in files named POSCAR1, POSCAR2, ... ; FORCEDISP1, FORCEDISP2, .... The FOCEX code checks consistency between the atomic structures of the supercell and the primitive cell described in structure.params. Force data may come from supercells of different shapes. Data for each shape is stored in a different input POSCARi, ... ; FORCEDISPi, ... i=1,2,… . Typically one large supercell can be used to extract the harmonic FCs, which can be of long range, and a smaller supercell can be used for large-amplitude anharmonic displacements. The files POSCARi contain the equilibrium or reference coordinates of the atoms in the supercell number i, in the POSCAR format defined in VASP.
@@ 3.10 Input files
3) The information regarding the reference supercells are read from files POSCAR1, POSCAR2, ... written in the same format as VASP POSCAR file. The corresponding cartesian displacements (in Ang) and forces (in eV/Ang) are read from files FORCEDISP1, FORCEDISP2, .... Each FORCEDISPi file contains all the force-displacement snapshots corresponding to the $i^{th}$ supercell defined in POSCARi. Each snapshot is separated from the next one by two lines, the first of which should contain the word ”POSITION” and the second one contains the total energy of that snapshot. The N following lines contain the cartesian coordinates of the N atoms followed by the forces acting on them (6 columns in total).
4) The file kpbs.params contains the list of k-points to use for the phonon band structure calculations. After the force constants are extracted, the FOCEX code also computes the harmonic properties, namely the phonon band structure, elastic and thermodynamic properties.
@@ 3.11 Output files
4) log-<extra>.dat contains the log of the run, where the parameters ”¡extra¿” label the type of the run, as described in the file ios.f90.
@@
9) ibz_dos.dat contains the Gaussian-smeared DOS from k-points generated in the IBZ, while fbz_dos.dat contains the DOS generated using the tetrahedron method.
@@ 3.12 Utilities
Some utilities (read_outcar.f90 and read_qe.f90) are also provided to convert the log or output files from a single snapshot of VASP (OUTCAR) or Quantum Espresso to FORCEDISP format. A shell script xtract.sh then calls these executables, and appends the output from each snapshot into the file FORCEDISP1. In addition, files for plotting the outputs using the Gnuplot software are also provided. all.plt or bvel.plt plot the phonon band structure, Gruneisen parameters DOS, and the group velocities. KTICS.BS should be imported in this file. fcs.plt plots the harmonic force constants versus pair distance and visualizes them in the Wigner-Seitz cell associated with the supercell. qha.plt plots the results from the QHA calculations: total energy, free energy, strain, volume, heat capacity, entropy and bulk modulus versus temperature. kp.plt visualizes the full zone, irreducible zone and band structure k-points in 3D reciprocal space.
@@ 4 ANFOMOD
To update atomic positions, the standard velocity Verlet [23] algorithm is used. If canonical ensemble needs to be simulated, a Langevin thermostat, in which the damping rate can be adjusted, has been adopted. A good initial guess for the latter would be $1/(1000dt)$ where $dt$ is the MD time step.
@@ 6 Installation
The alatdyn software package is available on GitHub under GPL-3.0 license at https://github.com/KeivanS/Anharmonic-lattice-dynamics, where each code is stored in the corresponding subdirectory. A Makefile template for each code has been provided. After cloning the repository, the user can type make to generate all needed executables, or alternatively go into each subdirectory, edit the Makefile according to their system and needs, and run make. The codes can be compiled using the GNU Fortran Compiler gfortran or the Intel Fortran Compiler ifort, and are tested to be compatible with gfortran versions 7-11 and ifort versions 18-19. (Newer and/or older versions are likely to be compatible as well.) The only external library is the zhegv diagonalization routine from LAPACK. A successful compilation results in the creation of the alatdyn_BIN directory containing all the executables.
In a typical high-performance computing cluster environment, one should be able to compile the codes using commands similar to the following. For example, to compile FOCEX with GCC (assuming gfortran is the install Fortran compiler):
@@ 7 Example of Ge, and some tips
Here we show the example of Germanium as a system ran with FOCEX. We must emphasize that the force-displacements inputs stored in FORCEDISP1 are up to the user. For a choice of the supercell and magnitude of displacements, the best set of FCs will be extracted. But the results will change with a change in the supercell, magnitude of displacements, choice of the exchange-correlation potential, lattice parameter, chosen rank and chosen range of FCs. So it is up to the user to make sure the obtained results are converged. As a rule of thumb, a supercell of 100 atoms or more is usually enough to provide a very reasonable set of force constants. To get accurate harmonic force constants, displacements need to be small enough. Typically displacements on the order of or less than 0.05 Å  are reasonable. While many codes require moving one or two atoms at the most, in FOCEX all atoms maybe simultaneously moved, providing more information on FCs for the same data size. A reasonable scheme would also be to move all atoms as a superposition of the normal modes sampled from the canonical ensemble, if effective force constants at higher temperatures need to the be extracted. If normal modes are not known, as is usually the case, one can still adopt normal modes generated with an arbitrary but cheap force field in order to avoid two separate supercell DFT calculations. The rationale for this approach is that direction of normal modes is defined by symmetry and the environment of the atom and is less dependent on the details of the force field. Generation of such displacements sampled from the canonical ensemble is done in the utility file sc_snaps.x which generates supercell snapshots of a required number if a temperature of interest is given. In principle, one or two snapshots in which all atoms are moved maybe sufficient to generate harmonic force constants. If only harmonic FCs are required, it is best to generate snapshots in which only inequivalent atoms are moved. More snapshots will be needed if anharmonic FCs are also required, typically on the order of 20, depending on their rank and chosen range. The number of generated FCs is a polynomial function of the range and an exponential function of the rank. Therefore more data (snapshots) would be needed to solve for all the unknowns if either range or rank is increased. The FOCEX code prints out for a chosen rank and range how many unknown FCs will be generated and therefore at least as many equations on forces vs displacements (snapshots) need to be generated. Checks must be performed to make sure the generated FCs don’t change when the range is increased; just the furthest ones are added and are smaller than the existing ones. Another check for convergence is that the physical quantities of interest such as the phonon dispersion and the Gruneisen parameters do not appreciably change as ranges are increased. Largest (infinite norm) and average (first norm) error, as well as the percentage deviation (second norm) of the fitted forces are written at the end of the log file.
@@
For Ge, we have used a supercell of 216 atoms with 36 snapshots in which only one or two atoms have been moved. Forces were calculated using VASP[30, 31, 32, 33] with the GGA-PBEsol functional, at cutoff energy of 225 eV and a Gamma-centered 7x7x7 k-point mesh within the 3x3x3 supercell. The lattice constant of $a=5.702$was found using 14x14x14 k-mesh for the primitive cell optimization calculations. This is only 0.7% larger than the experimental value of $5.66$. All harmonic FCs within the supercell were included in the fitting as shown in Fig. 2. Cubic force constants were included up to the 6th nearest neighbors. The results on the phonon dispersion and Gruneisen parameters are displayed in Fig. 3. The slight discrepancy between the computed and experimentally measured[34] frequencies is most likely due to the larger PBE lattice constant of 5.7 Å, which is slightly larger than the experimental value of 5.66 Å. This is a 2.1% larger volume. Therefore for a Gruneisen parameter of +1 for instance, a 2.1% reduction in the phonon frequency should be expected. The Gruneisen parameters in Fig.3 show a very decent agreement with the values of Olego and Cardona [35].
@@
Figure 2: Output of the fcs.plt: On the left are the fitted Ge harmonic force constants in units of the conventional lattice parameter $a_{0}=5.7\AA$, when only up to 6th nearest neighbors are included in the fitting (plus sign) and when all pairs within the supercell are included (open circles). Right: To visualize included force constants, we show the Wigner-Seitz (WS) cell of the supercell (in blue) from the top, centered on an atom, the grid of primitive translation vectors (green spheres), and selected force constants (lines) connecting the central atom to its neighbors within the supercell. Top right: only up to 6th neighbors, bottom right: all FCs within the supercell. This can be visualized for every atom in the primitive cell, but for Ge both atoms are equivalent.
@@
Figure 3: Mode Gruneisen parameters in the irreducible Brillouin zone (left), phonon density of states (center), and dispersion of Ge (right) calculated by FOCEX. Large red dots are experimental data. Small dots in black show the phonon dispersion curve as a result of including only 6 nearest neighbors in the fitting. TA modes are very sensitive to the number of neighbors. Near the X point TA dispersion should be flat. This can be achieved if at least 9 nearest neighbors are included. Color bar shows the group velocity magnitude in km/s
@@
LD model 121 69 54.0 86.4 73 105 54 41 0.36 0.29
Ioffe 126 44 67.7 71.3 103 41 0.26
McSkimin 129 48.3 67.25 77 102.5 67.1 0.26
MP (DFT) 102 36 56 58 83.3 47 0.25
Table 1: Elastic constants of Ge crystal in GPa. Experimental data from McSkimin[36] is at T=298K, and those of Ioffe Institute [37] are at 300K. The reported value of 0.19 on MP is not consistent with $-S_{12}/S_{11}=0.25$, but we are reporting the latter. For the Young and shear moduli and the Poisson ratio, we are reporting them for both deformations along [100] and the average value denoted by the subscript iso.
Group velocities are also compared to the data in Ioffe site[37]. Results, which are presented in table 2, show relatively good agreement. It is worth noting that there is a simple relationship between group velocities near the zone center (sound speeds) and elastic constants in a cubic crystal. All 3 elastic constants can be extracted from the 3 speeds of sound along [110]:
$\rho\,v_{L}^{2}(100)=C_{11}\,\,\rho\,v_{T1}^{2}(100)=\,\rho\,v_{T2}^{2}(100)=C_{44}\\ \rho\,v_{L}^{2}(111)=\left(\frac{C_{11}+2C_{12}+4C_{44}}{3}\right)\,\,\rho\,v_{T1}^{2}(111)=\rho\,v_{T2}^{2}(111)=\left(\frac{C_{11}-C_{12}+C_{44}}{3}\right)\\ \rho\,v_{L}^{2}(110)=\left(\frac{C_{11}+C_{12}}{2}+C_{44}\right)\,\,\rho\,v_{T1}^{2}(110)=C_{44}\,\,\rho\,v_{T2}^{2}(110)=\left(\frac{C_{11}-C_{12}}{2}\right)$ (31)
@@
Table 2: Group velocities (km/s) of the three acoustic branches compared to experimental data from McSkimin[36] and Ioffe Institute[37] taken at room temperature. The largest number corresponds to the longitudinal mode. Along [100] and [111] the two transverse acoustic branches are degenerate.
Figure 4: Left: Gruneisen parameter and linear thermal expansion versus T. Center: Total energy, free energy, specific heat and entropy as a function of T. Right: strain or volume expansion versus T.
Thermodynamic properties within QHA can also easily be computed. This requires the knowledge of Gruneisen parameters in order to predict phonon dispersion at different volumes. To find the equilibrium volume at a given temperature, the harmonic free energy is then minimized with respect to the volume. Figure 4 shows the plots of thermodynamic potential and its temperature derivatives, as well as the Gruneisen parameter and the linear thermal expansion. Although the zero CTE point at 50K [37] is correctly predicted, the CTE value is underestimated by QHA. Experimental data is $610^{-6}$ at 500K. There is also a measurement of CTE by Sparks[38], which shows a minimum at 30 K. These two low-temperature data points are shown in Fig. 4.
Finally, we must add that this code has been used many times and its results validated in the past 15 years or so. The results on force constants or phonons and lattice thermal conductivity have been published as follows: Silicon[39, 40, 41, 42], half-heuslers[43] FeSb2[44], graphene[45, 46], GaAs[47], PbSe and PbTe[48, 49], Bi and Sb[46].
@@
The Anharmonic LAttice DYNamics (alatdyn) suites of codes may be accessed and downloaded from
@@
https://alatdyn.readthedocs.io/en/latest/index.html